OpenStudy (anonymous):

Prove: (cosecx-secx)/(cosexc+secx)=(1-tanx)/(1+tanx)

OpenStudy (harsimran_hs4):

lets begin can you change the LHS in terms of sin and cos only

OpenStudy (anonymous):

\[\begin{align*}\frac{\csc x-\sec x}{\csc x+\sec x}&=\frac{1-\tan x}{1+tan x}\\\\ \frac{\frac{1}{\sin x}-\frac{1}{\cos x}}{\frac{1}{\sin x}+\frac{1}{\cos x}}&=\\\\ \frac{\frac{\cos x-\sin x}{\sin x\cos x}}{\frac{\cos x+\sin x}{\sin x\cos x}}&=\\\\ \frac{\cos x-\sin x}{\cos x+\sin x}&=\\\\ \frac{\cos x-\sin x}{\cos x+\sin x}\cdot\frac{\frac{1}{\cos x}}{\frac{1}{\cos x}} &=\\ \vdots&=\end{align*}\]

OpenStudy (anonymous):

Thank you! ^_^